OpenStudy (anonymous):

Solve for x. 20x - 2 = 18x - 4

OpenStudy (anonymous):

\[ 20x - 2 = 18x - 4 \] subtract \(18x\) form both sides add \(2\) to both sides divide both sides by \(2\)

OpenStudy (anonymous):

step: collect like terms and solve.

OpenStudy (anonymous):

20x - 2 = 18x - 4 20x = 18x - 2 2x = - 2 x = -1

OpenStudy (anonymous):

i did all the steps